OpenStudy (anonymous):

[6.02] and [6.03] Part 1: Solve the following system of equations. Please show your work to receive full credit. x - y = 10 2x + y = 2 Part 2: Explain which method you chose to solve the system and why you felt it was the best choice.

OpenStudy (anonymous):

If anyone knows, I will fan them, and give them a medal. :)

OpenStudy (anonymous):

Use the "Elimination" method. Your equations as they are are suitable for this method. x - y = 10 2x + y = 2 Notice that if we "combine" the quations, the Y variable will yield ZERO. (-Y+Y=0). Now we "combine the other terms, namely 3x+0y=12 or 3x=12 Therefore x=4 Now substitute this value into EITHER equation to solve for Y. I will choose the first equation. 4-y=10 or -y=6 Therefore y=-6. Hope that helps.

OpenStudy (anonymous):

Thats not right

OpenStudy (anonymous):

LOL, thx! I needed that! do you know inequalities? Wait. what is it then?

OpenStudy (anonymous):

first you have to change the equation to: x=y+13 then plug them in so you'd write it.... 2(y+13)+y=2 2y+26+y=2 3y+26=2 3y=-24 y=-8

OpenStudy (anonymous):

So it is the substitution method?

OpenStudy (anonymous):

after you plug it into the first equation you get x=7

OpenStudy (anonymous):

yes it is the substitution method!

OpenStudy (anonymous):

Ok, thx! Do you know this one?

OpenStudy (anonymous):

Hmmm?

OpenStudy (anonymous):

6.05] Two cars leave town going in opposite directions. One car is traveling 60 miles per hour (mph), and the other car is traveling 40 mph. How long will it take before the cars are 300 miles apart? 2 hours 3 hours 4 hours 5 hours

OpenStudy (anonymous):

iNeedTheHelp, check out the Wolfram solution (below). http://www.wolframalpha.com/input/?i=x-y%3D10%2C+2x%2By%3D2 My solution is correct.

OpenStudy (anonymous):

@onmng314 Use my solution and you will see that mine is true plug in the numbers using the substitution method and @MiaKitty the answer is b (: I have to go now good luck on your test fan me and ask me stuff if you have any more questions

OpenStudy (anonymous):

Ok, thx for your help! :D

OpenStudy (anonymous):

All I've done is given you another way to verify the solution to your first problem. iNeedTheHelp simply asked you to take his word for his solution. Check out the graphical interpretation I linked using Wolfram, and it is hard to argue with the conclusions. Good luck.

OpenStudy (anonymous):

Ok, thx! I will. :)

OpenStudy (anonymous):

Then give me the medal :)

OpenStudy (anonymous):

I don't know which is right, everyone says the website you used is untrustworthy, so I don't know. :(

OpenStudy (anonymous):

Wolfram, really? It is probably the ONLY source to check the answers to calculus problems. My classmates and I routinely use Wolfram to verify solutions. But....I won't try to convince you ;)

OpenStudy (anonymous):

Probs for calculus you know, but Algeba, most likely not,. :( I hate inequalities!

OpenStudy (anonymous):

You could even verify the system of equations problem using your graphing calculator. They are simply two lines. Either they intersect or they don't. If they do, it will be at a point (X,Y), which I have given. If you can, enter the equations and use the INTERSECT option to get the solution. So much work, I know :)

OpenStudy (anonymous):

yes, anyway, thx for your help! I have to check this out. :)

OpenStudy (anonymous):

PS: Make sure to rearrange your equations so that y=(everything else). That is how your calculator recognizes functions. x - y = 10 2x + y = 2 is the same as y=x-10 y=-2x+2 Use these when you input into your calc. Byeeeeeeeeeeeeeee

OpenStudy (anonymous):

Byeee, :) I fanned you!